OpenStudy (anonymous):

A kipping robe under constant vibation has a 5.0cm amplitude and an 8.0 second period. Wat vertical position is the rope at when t = 5.0s? b) Find two times when that same rope's position is +2.0 cm. c) Draw 2 cycles of this wave

OpenStudy (rob1525):

So, we know that the rope is making a sin wave right with a amplitude of 5 and period of 8 sec. We know that the sin wave has a period of 2pi and aplitude of 1. You need to change the amplitude and period of the sine wave to simulate the rope, then plug in 5 to find the height. The form y= A*sin(x).

OpenStudy (anonymous):

what equation would I use?

OpenStudy (rob1525):

Well, the amp is 5 so 5*sin(x) and now for the period is changed by changing x. Thus x now pi*x/4

OpenStudy (anonymous):

how do we know what x is?

OpenStudy (rob1525):

plug pi*x/4 into x. so y= 5*sin(pi*x/4). and solve by putting 5 in for x to find the height.

OpenStudy (anonymous):

I got 1.25

OpenStudy (anonymous):

nvm I got 0.3423...is that right?

OpenStudy (anonymous):

Help please!

OpenStudy (rob1525):

i got -3.54. Is your calc in radians?

OpenStudy (anonymous):

no should it be?

OpenStudy (anonymous):

and how can it be negative? the vertical position can be negative?

OpenStudy (rob1525):

Yes, your calc needs to be in radians to get the position of height. When you graph a sin wave it makes a wave above and below the x-axis right? So even though the value you get is negative its realy just 3.54 units away from the x-axis. can you graph sin(x)?

OpenStudy (anonymous):

alright makes sense, how about B?

OpenStudy (rob1525):

okay, so the sin wave periodic which means that the function y= 5*sin(pi*x/4) repeats over and over again. So their asking when is the rope at a height of positive 2 (above the x axis).

OpenStudy (anonymous):

so I would put 2 in for x?

OpenStudy (rob1525):

no x is time not height. so when is time = 2

OpenStudy (anonymous):

you mean when is height = 2?

OpenStudy (rob1525):

basicly 2=y= 5*sin(pi*x/4) and solve for x or just graph it and trae when height is 2.

OpenStudy (anonymous):

How would I graph it ?

OpenStudy (rob1525):

if you have graphing calc in the y editor put 5*sin(pi*x/4).

OpenStudy (anonymous):

Would the answer to b) be approx 3.46?

OpenStudy (anonymous):

I plugged my answer in on http://www.meta-calculator.com/online/

OpenStudy (rob1525):

Ya. looks close to what I got. but there are 2 others.

OpenStudy (anonymous):

approx 0.528?

OpenStudy (anonymous):

and approx -4.5?

OpenStudy (rob1525):

i ment 1 other

OpenStudy (anonymous):

so 0.528?

OpenStudy (anonymous):

0.5568*?

OpenStudy (rob1525):

Yes .528 is close to what i got.

OpenStudy (anonymous):

how can I find those numbers out using equations?

OpenStudy (rob1525):

not the negative answer because the rope frequency is from 0 to 8 sec.

OpenStudy (anonymous):

ah okay

OpenStudy (anonymous):

how can i find that .528 and 3.46 with equations?

OpenStudy (rob1525):

let me see

OpenStudy (rob1525):

Okay, so you set y= 5*sin(pi*x/4) to 2. 2= 5*sin(pi*x/4). then divide 5 so 2/5=sin(pi*x/4). Now use arcsin on both sides to undo sin then isolate x by multipling by 4/pi on both sides.

OpenStudy (anonymous):

I get 30? :S

OpenStudy (anonymous):

arcsin0.4 = 23.46

OpenStudy (rob1525):

okay, x= sin^-1(2/5)*(4/pi) = .524

OpenStudy (anonymous):

I plugged exactly that in and i get 30.02

OpenStudy (anonymous):

oh i was in degrees that explains it

OpenStudy (anonymous):

how can I get the other spot?